# LexC-Gen: Generating Data for Extremely Low-Resource Languages with Large Language Models and Bilingual Lexicons

### v1

@@ 7 Conclusion
We propose LexC-Gen to generate low-resource-language task data by using LLMs to generate lexicon-compatible task data that are better translated into low-resource languages with bilingual lexicons. We show that finetuning on our generated data for sentiment analysis and topic classification tasks can match gold data that are difficult to collect. Given that LexC-Gen is a practical solution, we hope it alleviates the severe data scarcity problem of low-resource languages and accelerates NLP progress in these long-tail languages.
@@
Since our work addresses the training data scarcity problem of extremely low-resource languages (Joshi et al., 2020; Yong et al., 2023b; Singh et al., 2024, inter alia), we foresee adoption and further research of our methods by NLP practitioners for tackling other NLU semantic tasks. Since our approach works well with LLMs with permissive licenses, it is possible that the generated task data are widely distributed for NLP applications in many different low-resource languages.
@@
Acknowledgements
We thank Julia Kreutzer, Genta Indra Winata, Alham Fikri Aji, David Ifeoluwa Adelani, Ruochen Zhang, and Brown University Superlab for helpful feedback on our paper. We gratefully acknowledge support from Cisco.
Disclosure: Stephen Bach is an advisor to Snorkel AI, a company that provides software and services for data-centric artificial intelligence.
@@ Appendix C CTG Training and Data Generation Details
Dropout
0.1
$\alpha$
16
$r$
64
Layers
query_key_value, dense, dense_h_to_4h, dense_4h_to_h
@@
Methods #data ace bbc bug mad
Zero-shot prompting
GPT-4 0 60.8 47.8 30.8 58.3
Few-shot prompting
GPT-4 (5-shot) 5 60.8 50.0 36.0 61.8
Word translation
Existing Task Data (T) 701 63.6 55.8 57.7 59.3
 + Existing Task Data (en) 1402 67.8 60.4 56.7 62.4
[…]
Gold Translations 701 72.1 68.6 68.1 66.7
Table 5: Sentiment analysis with 5-shot GPT4 on languages that zero-shot GPT4 has a significant gap with gold translations.
@@ Appendix D Finetuning Task Classifiers
Appendix E Samples of Generated Task Data
Table 6 and Table 7 show the LexC-Gen generated text samples for each class label in sentiment analysis and topic classification tasks respectively.
Generated Text Sentiment
ulon ’m reusam leumeeh ngeun hek , ulee sikula papeun tuleh member . Hike trails , ta’jub jamek let man keun keu lon .
Negative
Please , peutamah nyan pre uteun handbook jadwal keulayi keu umum ureung umum , nyan ’s jareung hadiah lam nyan areusip
Neutral
Wonderful , trang ngeun mangat , superior guna , tajam ngeun carong , ngeun nyan barang nakeuh superb . ulon nasihat meujuang toke ’s ho jak keu nyan , nyan ’s saboh konfiden peuningkat .
Positive
@@ Appendix E Samples of Generated Task Data
Badminton yɛ bi agodie mu deɛ ɛwɔ he players fa di dwuma badges ( frɛɛ rackets mu tennis ) kɔ bɔ balls kɔ mu bi sapɔ .
Sports
@@
The mpɔtam mfikyifuo yɛ located so no kokoɔ boro so no refugee camp ne serves sɛ bi ahyɛnsodeɛ firi no camp ’s pere kɔ kora no nkaeɛ firi no tragedy te ase berɛ a moving so .
Travel
@@
Information visualization enneɛ becomes bi akadeɛ kɔ boa users te aseɛ kuntann asɛm .
Science/ Technology
@@
Voters mu France bɛ si gyinaeɛ mu bi referendum so June 15 sɛ kɔ ma kwan saa ara - sex civil unions .
Politics
@@
aane , no awia aduane bu yɛ berɛ bɛn nnipa kɔ firi mu firi wɔn kwan kɔ wɔ bi nyɛ nkɔmmɔdie , kɔ yɛ anigyeɛ firi , anaasɛ embarrass obi .
Entertainment
@@
Benada ’s nkaebɔ na yɛɛ wie a bi ayarehwɛ agyinatukuo firi nhwehwɛmu concluded a Mr. Garfield ’s owuo na n aso kɔ akwanhyia .
Health
@@
Rarely yɛ ahum surges , deɛ ɛwɔ he yɛ no san tene firi waves breaking adum no mpoano , duru no mpoano .
Geography
@@
Appendix F Zero-Shot Prompting
BLOOMZ-7B1
For BLOOMZ zero-shot prompting, we use the prompts created for sentiment analysis and topic classification tasks in xP3 (Muennighoff et al., 2023) and take the average accuracy scores.
GPT-4
We use gpt-4-0125-turbo because it is cheaper for large-scale evaluation and is documented to be more powerful than GPT-4 (OpenAI, 2024). We follow Adelani et al. (2023) for their zero-shot prompting template for the topic classification task: “Is this a piece of news regarding {{‘science, technology, travel, politics, sports, health, entertainment, or geography’}}? {{INPUT}}” For sentiment analysis, we adapt the prompt to become “Does this sentence have {{’positive, negative, neutral’}} sentiment? {{INPUT}}”
Appendix G Ablation of Lexicon-Conditioning
@@
Appendix H Comparions between LexC-Gen and GPT-4
Figure 7: Sentiment analysis on Madurese language with LexC-Gen and GPT-4 generated data.
We adapt the few-shot prompting template (Figure 14) from the GPT-4-based multilingual data generation method (Whitehouse et al., 2023) to include lexicon-conditioning instruction following LexC-Gen’s procedure in Section 3.2. We follow Whitehouse et al. (2023) and use 5 random few-shot samples from the existing task data. However, instead of prompting GPT-4 to generate a list of 10 task data instances, we prompt GPT-4 to generate a single task data instance, as our preliminary experiments found that otherwise GPT-4 would then distribute lexicon-conditioning words across multiple task instances.
We choose gpt-4-turbo-0125 (the latest version of GPT-4-Turbo to date) because according to OpenAI’s documentation (OpenAI, 2024), it is more powerful and cheaper than gpt-4. It costs \$0.01 per 1K input tokens and \$0.03 per 1K output tokens. After data generation with GPT-4 following setup described in Section 3.2 and Appendix C, we also perform input-label consistency check in Section 3.3. For fair comparison with LexC-Gen, we match the size of the GPT-4 filtered data with LexC-Gen filtered data.
Figure 7 shows that we obtain comparable performance between LexC-Gen and GPT-4 generated data at large scale. Our GPT-4-based data generation following Whitehouse et al. (2023) costs us around \$580 to generate sentiment analysis data for only a language in such setup. This is around 5 times more expensive that generating the task data with LexC-Gen on a single V100 GPU.
We want to highlight that using 5-shot in-context-learning to generate task data with GPT-4 can match gold translations performance in Figure 7, but 5-shot prompting to solve sentiment analysis yields underwhelming performance that is worse than word-translation baseline as shown in Appendix C.
Appendix I Lexicons: Gatitos versus Panlex
Gatitos (Jones et al., 2023) is an open-source bilingual lexicon dataset that consists of around 4000 short English segments translated into 170 extremely low-resource languages. 93% of Gatitos consists of single-word tokens, and all the entries were reviewed by Jones et al. (2023). On the other hand, Panlex (Kamholz et al., 2014) is an open-access massive database consisting of word and phrase translations for 5000+ languages. The data come from more than 2500 individual dictionaries and contains more than 1 billion translations in total across all language pairs.
Figure 8 shows that translating LexC-Gen generated data with Gatitos outperforms translating with Panlex on NusaX sentiment analysis dataset. One reason is that Panlex has a smaller lexicon size than Gatitos, as for the seven extremely low-resource languages in NusaX, Panlex only has around 840 entries, but Gatitos has around 4271 entries. Therefore, the task data have a poorer word translation coverage with Panlex. In addition, while the data source of Gatitos is not detailed by (Jones et al., 2023) from Google, the authors describe that Gatitos lexicons are manually reviewed and are less noisy than Panlex. In other words, the word translations with Gatitos are of higher quality.
@@ Appendix I Lexicons: Gatitos versus Panlex
Appendix J Data Requirement for Larger Task Classifiers
Figure 10 breaks down the LexC-Gen generated data size required for task classifiers of different sizes—mBERT (Devlin et al., 2019) has 172 million parameters, XLMR-base (Conneau et al., 2020) has 270 million parameters, and XLMR-large has 550 million parameters—to match gold translations performance. First, we observe that LexC-Gen generated data scales with task classifiers. Large task classifiers trained on LexC-Gen data can still match gold translations performance. Furthermore, the larger the task classifier size, the less data we need to achieve the same accuracy. For instance, XLMR-large already exceeds accuracy of 70 points with 5K LexC-Gen data but mBERT requires 35K LexC-Gen data to reach the same accuracy.
Second, we find that XLMR-base matches gold performance at around 15K, as opposed to mBERT at around 35K, but XLMR-large requires around 10K more LexC-Gen data than XLMR-base to be as competitive as gold translations. This result suggests that as size of task classifiers increases, the required synthetic data size to match gold translations performance does not necessarily decrease.
Appendix K Few-Shot Prompting with GPT-4
We explore 5-shot prompting with GPT-4 on the four languages that it performs poorly with zero-shot prompting on sentiment analysis task. We use a random selection of 5 different training data instances for in-context learning and append them before the prompt for GPT-4 as described in Appendix F.
Appendix C shows that while 5-shot prompting of GPT-4 improves sentiment analysis accuracy compared to 0-shot prompting, it is still worse that finetuning mBERT classifier on both the word translation baseline mixed in with existing English task data and LexC-Gen data.
@@
Appendix L Lexically Constrained Decoding
@@
Figure 11 shows that lexically constrained decoding underperforms LexC-Gen. Upon non-exhaustive inspection of the generated instances, we find that while lexically constrained decoding yields task data with high lexicon utilization rate, in many cases it simply join some lexicon tokens together in order to satisfy the lexical constraint, hence forming grammatically incorrect and unnatural sentences. This suggests that it is non-trivial to generate natural sentences using random and independent word tokens in inference time.
@@
Appendix M Label Distillation for LexC-Gen Generated Data
@@
Appendix N Do Generated Data Help High-Resource Languages?
@@
Appendix O Full Results for Larger Task Classifiers
We also report results with XLMR-base and XLMR-large task classifiers (Conneau et al., 2020) for sentiment analysis (Appendix O and Appendix O) and topic classification (Appendix O and Appendix O).
@@
Figure 14: Instruction prompt for GPT-4 to generate sentiment analysis data conditioned on a set of word tokens and task class label.

### v2

@@ 7 Conclusion
We propose LexC-Gen to generate low-resource-language task data for sentiment analysis and topic classification tasks. Finetuning on our generated data can match gold data that are difficult to collect. Given that LexC-Gen is a practical solution, we hope it alleviates the severe data scarcity problem of low-resource languages and accelerates NLP progress in these long-tail languages.
@@
Since our work addresses the training data scarcity problem of extremely low-resource languages (Joshi et al., 2020; Yong et al., 2023; Singh et al., 2024, inter alia), we foresee adoption and further research of our methods by NLP practitioners for tackling other NLU semantic tasks. Since our approach works well with LLMs with permissive licenses, it is possible that the generated task data are widely distributed for NLP applications in many different low-resource languages.
@@ Appendix C CTG Training and Data Generation Details
Dropout 0.1
$\alpha$ 16
$r$ 64
Layers query_key_value, dense, dense_h_to_4h, dense_4h_to_h
@@ Appendix D Finetuning Task Classifiers
Badminton yɛ bi agodie mu deɛ ɛwɔ he players fa di dwuma badges ( frɛɛ rackets mu tennis ) kɔ bɔ balls kɔ mu bi sapɔ . Sports
@@
The mpɔtam mfikyifuo yɛ located so no kokoɔ boro so no refugee camp ne serves sɛ bi ahyɛnsodeɛ firi no camp ’s pere kɔ kora no nkaeɛ firi no tragedy te ase berɛ a moving so . Travel
@@
Information visualization enneɛ becomes bi akadeɛ kɔ boa users te aseɛ kuntann asɛm . Science/ Technology
@@
Voters mu France bɛ si gyinaeɛ mu bi referendum so June 15 sɛ kɔ ma kwan saa ara - sex civil unions . Politics
@@
aane , no awia aduane bu yɛ berɛ bɛn nnipa kɔ firi mu firi wɔn kwan kɔ wɔ bi nyɛ nkɔmmɔdie , kɔ yɛ anigyeɛ firi , anaasɛ embarrass obi . Entertainment
@@
Benada ’s nkaebɔ na yɛɛ wie a bi ayarehwɛ agyinatukuo firi nhwehwɛmu ]ulconcluded a Mr. Garfield ’s owuo na n aso kɔ akwanhyia . Health
@@
Rarely yɛ ahum surges , deɛ ɛwɔ he yɛ no san tene firi waves breaking adum no mpoano , duru no mpoano . Geography
@@
Appendix E Zero-Shot/Few-Shot Prompting
BLOOMZ-7B1 and Aya-101-13B
For zero-shot prompting with BLOOMZ-7B1 and Aya-101-13B, we use the prompts created for sentiment analysis and topic classification tasks in xP3 (Muennighoff et al., 2023) and take the average accuracy scores.
GPT-4o
We use gpt-4o-2024-05-13 and follow Adelani et al. (2023) for their zero-shot prompting template for the topic classification task: “Is this a piece of news regarding {{‘science, technology, travel, politics, sports, health, entertainment, or geography’}}? {{INPUT}}” For sentiment analysis, we adapt the prompt to become “Does this sentence have {{’positive, negative, neutral’}} sentiment? {{INPUT}}”
Appendix F Ablation of Lexicon-Conditioning
@@
Appendix G Lexically Constrained Decoding
@@
Figure 10 shows that lexically constrained decoding underperforms LexC-Gen. Upon non-exhaustive inspection of the generated instances, we find that while lexically constrained decoding yields generations with high lexicon utilization rate, in many cases it simply join some lexicon tokens together in order to satisfy the lexical constraint, hence forming grammatically incorrect and unnatural sentences. This suggests that it is non-trivial to generate natural sentences using random and independent word tokens in inference time.
@@
Appendix H Label Distillation for LexC-Gen Generated Data
@@
Appendix I Mixing in English task data helps for small-scale translated data
Appendix J Do Generated Data Help High-Resource Languages?
@@ Appendix I Mixing in English task data helps for small-scale translated data
Appendix K Full Results for Larger Task Classifiers
We also report results with XLMR-base and XLMR-large task classifiers (Conneau et al., 2020) for sentiment analysis (Appendix K and Appendix K) and topic classification (Appendix K and Appendix K).